# Efficient Attention Mechanisms for Large Language Models: A Survey

### v2

@@ 3.3 Clustering Attention
RetrievalAttention [30] employs Approximate Nearest Neighbor Search (ANNS) for selecting critical K clusters. To address the challenge of the out-of-distribution nature between query and key vectors in the attention mechanism, it introduces an attention-aware vector search algorithm that adapts to the distribution of query vectors.
ClusterKV [31] select tokens at the granularity of semantic clusters, overcoming the issue of internal fragmentation of page-level retrieval methods such as Quest. After prefilling stage, tokens are clustered through K-means algorithm. The semantic similarity between token $i$ and $j$ are measured through the cosine similarity of key vectors $\mathcal{D}(i,j)=1-\frac{<k_{i},k_{j}>}{|k_{i}|\cdot|k_{j}}$. Semantic clusters are represented by their centroids $\mu_{1},\mu_{2},...,\mu_{C}\in\mathcal{R}^{d}$. At each decoding step, clusters are selected based on the ranking of query token $q$ and centroids $\mu_{i}$’s attention weights, i.e. $q\mu_{i}^{T}$.
MagicPIG [7] leverages Locality Sensitive Hashing (LSH) sampling to efficiently approximate attention computation. It employs LSH to map similar query and key vectors to the same hash buckets and offloads storage and partial computation to the CPU to address the KV cache bottleneck. It also introduces Oracle Top-K sampling as a better strategy than brute force Top-K.
@@ 3.4 Bidirectional Sparse Attention
Block sparsity is widely used in bidirectional sparse attention. BigBird [73] uses block-wise random attention, which acts as bridges to shorten indirect paths between tokens. Longformer [4] uses static global-local hybrid attention. It also relies on block-level sparsity with additional global and random links, facilitating structured computation and memory-efficient parallelism.
Clustering-based methods are also used in bidirectional sparse attention. Reformer [24] uses Locality-Sensitive Hashing (LSH) to assign similar tokens to the same bucket. Routing Transformer [48] performs online $k$-means clustering per layer. ClusterFormer [61] introduces a differentiable clustering module co-trained with downstream objectives. These methods reduce computation by grouping related tokens while maintaining performance through learned adaptability.
@@ 4.1 Pretrained Models with Uniform Efficient Attention
The RWKV project represents a sustained and influential effort to create a scalable Recurrent Neural Network (RNN) architecture that combines the parallelizable training of Transformers with the efficient inference of traditional RNNs [37]. For instance, the EAGLE (RWKV-5) series introduced matrix-valued states to increase capacity, while subsequent iterations like Finch (RWKV-6) [38] and Goose (RWKV-7) [39] incorporated dynamic recurrence and expressive state evolution mechanisms (e.g., a delta-rule) to enable more complex, data-dependent state transitions.
@@
The success of the Mamba architecture [17], with its data-dependent selection mechanism (Section 2.2.2), has spurred a wave of adoption and scaling initiatives from major research labs. Falcon Mamba [78] is based on the pure Mamba-based architecture to demonstrate performance competitive with leading Transformer models on a wide range of general-purpose language benchmarks, validating the architecture’s viability for such tasks while retaining its signature constant-time inference. Further evidence of this paradigm’s potential is provided by Codestral Mamba [35], built on the Mamba-2 architecture. While specialized for code generation, it achieves state-of-the-art results on relevant benchmarks and supports a 256k token context, demonstrating the scalability and effectiveness of the SSM approach within a complex and structured domain.
@@
MiniCPM-4 [57] introduces a two-stage sparse attention mechanism that dynamically selects relevant key-value blocks for each query token based on semantic similarity. MiniCPM-4 leverages InfLLM-v2, a Block Sparse Attention variant to replace standard Attention mechanism. Moreover, a lightweight LogSumExp approximation enables efficient top-k selection, making the method scalable to extremely long sequences. Together, these techniques allow MiniCPM-4 to balance fine-grained contextual awareness with tractable memory and compute requirements, making it a strong candidate for long-context modeling.
@@ 4.2 Pretrained Models with Hybrid Efficient Attention
GPT-3 [5] integrates a hybrid attention mechanism by interleaving dense and locally banded sparse attention layers, inspired by the Sparse Transformer [8]. Dense attention provides full-context modeling, while sparse layers adopt fixed or strided patterns to reduce the number of attended tokens. This design enables GPT-3 to efficiently scale to large model sizes using a fixed context window of 2048 tokens, balancing modeling capacity and computational efficiency.
@@
Jamba [28] and MiniMax-01 [26] combine linear and full attention layers to achieve an efficient trade-off between throughput and expressiveness. MiniMax-01 employs Lightning Attention across most layers, inserting Softmax-based full attention every eight layers. Jamba adopts a similar ratio, inserting one Transformer layer into every eight-layer Mamba block. Both achieve faster decoding and improved long-sequence performance by limiting the use of computationally intensive full attention.
@@
Gemma 3 [56], Command A [66], and LLaMA-4-Maverick [32] alternate between local and global attention layers, with a shared design philosophy of using global layers sparsely, e.g., every 4–6 layers, to enhance efficiency. While local layers adopt sliding-window patterns, the key difference lies in position encoding strategies. Gemma 3 modulates RoPE base frequencies—assigning 10K for local and 1M for global layers—to better capture long-range dependencies. Command A and LLaMA-4-Maverick mixes RoPE-based local layers with full attention layers that omit positional embeddings entirely, allowing stronger long sequence performance.
@@
Character.AI [6] interleaves local attention with sliding windows and sparse global attention layers applied every six layers. Especially, they resue global attention layer’s key-value representations across multiple non-adjacent layers. This KV sharing mechanism enables efficient long-context processing with reduced memory and latency overhead.
YOCO [53] and Phi-4-mini-flash [47] adopt a dual-decoder architecture that separates the prefill and generation phases. The Self-Decoder utilizes linear attention mechanisms such as RetNet and Sliding-Window Attention for both prefill and generation, while the Cross-Decoder is activated only during generation. A single-layer global KV cache is used throughout, allowing linear-time prefill and efficient decoding with minimal GPU memory consumption.
@@ 5 Outlook
6 Acknowledgment
This work was supported in part by National Key Research and Development Program of China under Grant No. 2020YFA0804503, National Natural Science Foundation of China under Grant No. 62272264, and ByteDance Doubao Large Model Fund Project under Grant No. CT20240909109354.

### v3

@@ 3.3 Clustering Attention
RetrievalAttention [liu2024retrievalattentionacceleratinglongcontextllm] employs Approximate Nearest Neighbor Search (ANNS) for selecting critical K clusters. To address the challenge of the out-of-distribution nature between query and key vectors in the attention mechanism, it introduces an attention-aware vector search algorithm that adapts to the distribution of query vectors.
ClusterKV [liu2024clusterkvmanipulatingllmkv] select tokens at the granularity of semantic clusters, overcoming the issue of internal fragmentation of page-level retrieval methods such as Quest. After prefilling stage, tokens are clustered through K-means algorithm. The semantic similarity between token $i$ and $j$ are measured through the cosine similarity of key vectors $\mathcal{D}(i,j)=1-\frac{<k_{i},k_{j}>}{|k_{i}|\cdot|k_{j}}$. Semantic clusters are represented by their centroids $\mu_{1},\mu_{2},...,\mu_{C}\in\mathcal{R}^{d}$. At each decoding step, clusters are selected based on the ranking of query token $q$ and centroids $\mu_{i}$’s attention weights, i.e. $q\mu_{i}^{T}$.
MagicPIG [chen2024magicpiglshsamplingefficient] leverages Locality Sensitive Hashing (LSH) sampling to efficiently approximate attention computation. It employs LSH to map similar query and key vectors to the same hash buckets and offloads storage and partial computation to the CPU to address the KV cache bottleneck. It also introduces Oracle Top-K sampling as a better strategy than brute force Top-K.
@@ 3.4 Bidirectional Sparse Attention
Block sparsity is widely used in bidirectional sparse attention. BigBird [bigbird] uses block-wise random attention, which acts as bridges to shorten indirect paths between tokens. Longformer [longformer] uses static global-local hybrid attention. It also relies on block-level sparsity with additional global and random links, facilitating structured computation and memory-efficient parallelism.
Clustering-based methods are also used in bidirectional sparse attention. Reformer [reformer] uses Locality-Sensitive Hashing (LSH) to assign similar tokens to the same bucket. Routing Transformer [routing] performs online $k$-means clustering per layer. ClusterFormer [clusterformer] introduces a differentiable clustering module co-trained with downstream objectives. These methods reduce computation by grouping related tokens while maintaining performance through learned adaptability.
@@ 4.1 Pretrained Models with Uniform Efficient Attention
The RWKV project represents a sustained and influential effort to create a scalable Recurrent Neural Network (RNN) architecture that combines the parallelizable training of Transformers with the efficient inference of traditional RNNs [rwkv]. For instance, the EAGLE (RWKV-5) series introduced matrix-valued states to increase capacity, while subsequent iterations like Finch (RWKV-6) [rwkv56] and Goose (RWKV-7) [rwkv7] incorporated dynamic recurrence and expressive state evolution mechanisms (e.g., a delta-rule) to enable more complex, data-dependent state transitions.
@@
The success of the Mamba architecture [mamba], with its data-dependent selection mechanism (Section 2.2.2), has spurred a wave of adoption and scaling initiatives from major research labs. Falcon Mamba [falcon_mamba] is based on the pure Mamba-based architecture to demonstrate performance competitive with leading Transformer models on a wide range of general-purpose language benchmarks, validating the architecture’s viability for such tasks while retaining its signature constant-time inference. Further evidence of this paradigm’s potential is provided by Codestral Mamba [codestral_mamba], built on the Mamba-2 architecture. While specialized for code generation, it achieves state-of-the-art results on relevant benchmarks and supports a 256k token context, demonstrating the scalability and effectiveness of the SSM approach within a complex and structured domain.
@@
MiniCPM-4 [minicpm4] introduces a two-stage sparse attention mechanism that dynamically selects relevant key-value blocks for each query token based on semantic similarity. MiniCPM-4 leverages InfLLM-v2, a Block Sparse Attention variant to replace standard Attention mechanism. Moreover, a lightweight LogSumExp approximation enables efficient top-k selection, making the method scalable to extremely long sequences. Together, these techniques allow MiniCPM-4 to balance fine-grained contextual awareness with tractable memory and compute requirements, making it a strong candidate for long-context modeling.
@@ 4.2 Pretrained Models with Hybrid Efficient Attention
GPT-3 [gpt3] integrates a hybrid attention mechanism by interleaving dense and locally banded sparse attention layers, inspired by the Sparse Transformer [sparsetransformer]. Dense attention provides full-context modeling, while sparse layers adopt fixed or strided patterns to reduce the number of attended tokens. This design enables GPT-3 to efficiently scale to large model sizes using a fixed context window of 2048 tokens, balancing modeling capacity and computational efficiency.
@@
Jamba [jamba] and MiniMax-01 [minimax01] combine linear and full attention layers to achieve an efficient trade-off between throughput and expressiveness. MiniMax-01 employs Lightning Attention across most layers, inserting Softmax-based full attention every eight layers. Jamba adopts a similar ratio, inserting one Transformer layer into every eight-layer Mamba block. Both achieve faster decoding and improved long-sequence performance by limiting the use of computationally intensive full attention.
@@
Gemma 3 [gemma3], Command A [rnope], and LLaMA-4-Maverick [llama4] alternate between local and global attention layers, with a shared design philosophy of using global layers sparsely, e.g., every 4–6 layers, to enhance efficiency. While local layers adopt sliding-window patterns, the key difference lies in position encoding strategies. Gemma 3 modulates RoPE base frequencies—assigning 10K for local and 1M for global layers—to better capture long-range dependencies. Command A and LLaMA-4-Maverick mixes RoPE-based local layers with full attention layers that omit positional embeddings entirely, allowing stronger long sequence performance.
@@
Character.AI [characterai] interleaves local attention with sliding windows and sparse global attention layers applied every six layers. Especially, they resue global attention layer’s key-value representations across multiple non-adjacent layers. This KV sharing mechanism enables efficient long-context processing with reduced memory and latency overhead.
YOCO [yoco] and Phi-4-mini-flash [phi4miniflash] adopt a dual-decoder architecture that separates the prefill and generation phases. The Self-Decoder utilizes linear attention mechanisms such as RetNet and Sliding-Window Attention for both prefill and generation, while the Cross-Decoder is activated only during generation. A single-layer global KV cache is used throughout, allowing linear-time prefill and efficient decoding with minimal GPU memory consumption.
@@
5 TAKEAWAYS
Based on the current trajectory of both academic research and industrial deployment, several key consensus points emerge regarding the design of efficient attention mechanisms.
The Superiority of Hybridity
There is a growing consensus that while purely linear attention architectures offer impressive inference speeds, they frequently incur a non-trivial performance penalty in complex reasoning and ultra-long context retrieval. Consequently, hybrid models have emerged as the preferred architecture. These designs not only maintain the near-lossless performance of standard Transformers but demonstrate the potential to exceed them by balancing global associative memory with precise local retrieval.
The Gradient Bottleneck in Sparse Training
Training sparse attention models from scratch remains significantly more challenging than dense counterparts. This performance gap primarily stems from gradient sparsity, where restricted connectivity leads to suboptimal optimization dynamics and lower learning efficiency. While self-distillation has shown promise in bridging this gap, it remains an auxiliary solution, indicating that sparse-from-scratch strategies require further foundational innovation.
Complementary Strengths of Linear and Sparse Attention
From a unified functional perspective, linear and sparse mechanisms are increasingly viewed as complementary rather than competitive. In a modern efficient stack, linear attention serves as an effective state compressor for fast inference and KV cache reduction, while sparse attention acts as a high-fidelity module to replace full attention during decoding. A robust architecture leverages this synergy to navigate the Pareto frontier of computational throughput and predictive accuracy.
The Critical Role of Hardware-Algorithm Co-design
A recurring takeaway is that theoretical FLOPs reduction does not inherently translate to wall-clock speedups. Modern efficient attention is inextricably bound to hardware optimization. For instance, the utility of linear attention hinges on switching between recurrent and parallel forms based on the workload, while sparse attention’s success depends on resolving the tension between theoretical sparsity and the massive parallelism required by modern GPUs.